OpenStudy (anonymous):

could you help me solve for r when Cr^-2

OpenStudy (saifoo.khan):

Online College?

OpenStudy (anonymous):

yeah I am in online college But I am a little stuck on this

OpenStudy (saifoo.khan):

Alright, THis is the Apendix question. Just a sec.

OpenStudy (anonymous):

We only do these on Wednesdays between 1 and half past..

OpenStudy (saifoo.khan):

W = C/r^2 r^2=C/W r = sqrt(c/w)

OpenStudy (saifoo.khan):

lol @e-studier.

OpenStudy (anonymous):

thank you it made sense once i saw it